# Dynamic Scene Understanding through Object-Centric Voxelization and Neural Rendering

### v1

@@ 3.4 Training Procedure
DynaVol-S involves three stages in the training phase: First, a warmup stage that learns to obtain a basic understanding of the geometric, semantic, and dynamic priors. Second, an initialization stage for the voxelized occupancy probabilities $\mathcal{V}^{\text{M}}$, involving the connected components algorithm. Third, an end-to-end optimization stage that jointly optimizes $\{\mathcal{V}^{\text{D}},\mathcal{V}^{\text{C}},\mathcal{V}^{\text{M}},\{\mathbf{z}_{n}\}_{n=1}^{N},f_{\psi}\}$ to refine the voxelized object occupancy.
@@
20:     $\mathcal{V}^{\text{M}}[n,:,:,:]=\text{one-hot}(\text{cl}_{k})$triangleright Generate 4D voxel grids
@@
22:  Rendering 2D masks $\widehat{P}(\mathbf{r})$ with $\mathcal{V}^{\text{M}}$
@@ 3.4.1 Warmup
To reduce the difficulty of learning the object-centric 4D voxel grids $\mathcal{V}^{\text{M}}$, we first use $\{\mathbf{I}_{t}\}_{t=1}^{T}$ and $\{\mathbf{F}_{t}\}_{t=1}^{T}$ to warm-up the density grids $\mathcal{V}^{\text{D}}$, the color-related feature grids $\mathcal{V}^{\text{C}}$, the semantic feature grids $\mathcal{V}^{\text{S}}$, the slot features $\mathbf{S}$, and the dynamics module $f_{\psi}(\cdot)$. Additionally, we train an extra forward dynamics module $f^{\prime}_{\xi}(\cdot)$ and a non-compositional neural renderer $N_{\phi^{\prime}}(\cdot)$. The overall objective is defined as $\mathcal{L}_{\text{Warm}}=\mathcal{L}_{\text{Render}}+\alpha_{p}\mathcal{L}_{\text{Point}}+\alpha_{b}\mathcal{L}_{\text{Bg-Ent}}+\alpha_{c}\mathcal{L}_{\text{Cyc}}+\mathcal{L}_{\text{Recon}}+\alpha_{e}\mathcal{L}_{\text{Ent}}$. This framework divides its focus between scene reconstruction, covered by the first four terms, and the representation of semantic features, addressed by the last two terms. The hyperparameter values of scene reconstruction are adopted from prior literature [22].
At a specific timestamp, we first use the rendering loss $\mathcal{L}_{\text{Render}}$ between the predicted and observed pixel colors, the background entropy loss $\mathcal{L}_{\text{Bg-Ent}}$, and the per-point RGB loss $\mathcal{L}_{\text{Point}}$ following DVGO [23] as basic objective terms for neural rendering:
@@
where $\mathcal{R}$ is the set of sampled rays in a batch, $P$ is the number of sampling points along ray $r$, and $\widehat{w}_{l}^{r}$ is the color contribution of the last sampling point obtained by $\widehat{w}^{r}_{l}=T_{P}(1-\exp(-\sigma_{P}\delta_{P}))$. $\mathcal{L}_{\text{Bg-Ent}}$ can be viewed as a regularization to encourage the renderer to concentrate on either foreground or background. To enhance the coherence of the learned canonical-space dynamics, we train an additional module $f^{\prime}_{\xi}(\cdot)$ using the following cycle-consistency loss:
@@
where $x^{\prime}_{i}=x_{i}+f_{\psi}(x_{i},t),i\in[1,P]$. Furthermore, we incorporate the reconstruction loss $\mathcal{L}_{\text{Recon}}$ between the predicted and observed pixel semantic features, as well as the entropy loss $\mathcal{L}_{\text{Ent}}$ on the semantic probabilities to regularize them towards a one-hot distribution:
3.4.2 Object-Centric Voxel Grids Initialization
We initialize the 4D object-centric grids $\mathcal{V}^{\text{M}}$ with two steps: (1) the feature graph generation and connected components computation and (2) a post-processing step with Conditional Random Field (CRF). As illustrated in Alg. 1, the entire process takes as inputs the voxel position set $\{\mathbf{x}_{k}\}$ with a size of $N_{s}=N_{x}\times N_{y}\times N_{z}$, the voxel density set $\{\sigma_{k}\}$ from $\mathcal{V}^{\text{D}}$, the view-independent colors $\{\mathbf{c}_{k}^{\text{ind}}\}$ generated by $N_{\phi^{\prime}}$, and the corresponding semantic labels $\{y_{k}\}$ obtained by $\mathrm{argmax}_{N}\left(\mathcal{A}\right)$ in Eq. (5). We also consider the forward canonical-space transitions generated by $f^{\prime}_{\xi}(\cdot)$. Besides, we set five hyperparameters in advance, where $\delta_{\text{den}}$ represents the density value threshold, $\delta_{\text{pos}}$ represents the position distance threshold, $\delta_{\text{rgb}}$ represents the RGB distance threshold, $\delta_{\text{vel}}$ represents the velocity distance threshold, and $N$ represents the maximum number of objects.
Furthermore, as an optional post-processing step in DynaVol-S, we employ a Conditional Random Field (CRF) to refine $\mathcal{V}^{M}$, following the common practice in video object segmentation, this step is beneficial for mitigating potential noise introduced by the connected components algorithm, thereby smoothing the segmentation results. The CRF method leverages RGB information to refine 2D mask predictions. Although this step could potentially compromise the 4D consistency of segmentation results, we address this by using the refined masks as pseudo-labels in the subsequent stage. This allows us to finetune the 4D voxel grids, $\mathcal{V}^{\text{M}}$, thereby ensuring consistency in object-centric representations.
3.4.3 Multi-Grids Joint Optimization
In this stage, we finetune the dynamics module $f_{\psi}$, the geometric grids $\mathcal{V}^{\text{D}}$, the appearance grids $\mathcal{V}^{\text{C}}$, and the object-centric grids $\mathcal{V}^{\text{M}}$ obtained in the first two stages. We finetune the parameters of $(\mathcal{V}^{\text{D}},\mathcal{V}^{\text{C}},\mathcal{V}^{\text{M}},\psi)$ and begin training ($(\{\mathbf{z}_{n}\}_{n=1}^{N},\phi)$ from scratch. Following an end-to-end training scheme, the dynamics module, object-centric latent codes, and compositional renderer collaboratively contribute to object-centric dynamic scene understanding. The loss function in this stage is defined as $\mathcal{L}_{\text{Dyn}}=\mathcal{L}_{\text{Render}}+\alpha_{p}\mathcal{L}_{\text{Point}}+\alpha_{b}\mathcal{L}_{\text{Bg-Ent}}$, complemented by a cross-entropy loss between the rendered 2D segmentations $\widehat{P}(\mathbf{r})$ from $\mathcal{V}^{\text{M}}$ and the pseudo-labels $\bar{P}(\mathbf{r})$ derived from the second stage: $\mathcal{L}_{\text{CE}}=-\sum_{n=1}^{N}\bar{P}(\mathbf{r})_{n}\log\widehat{P}(\mathbf{r})_{n}$.
4 Experiments
@@ 4.1 Experimental Setup
Datasets. We build the $8$ synthetic dynamic scenes in Table II using the Kubric simulator [25]. Each scene spans $60$ timestamps and contains different numbers of objects in various colors, shapes, and textures. The objects have diverse motion patterns and initial velocities. All images have dimensions of $512\times 512$ pixels. We also adopt real-world scenes from HyperNeRF [8] and $\text{D}^{2}$eRF [9], as shown in Table III and Table V. For the synthetic scenes, we follow D-NeRF [17] to employ images collected at viewpoints randomly sampled on the upper hemisphere. In contrast, real-world scenes are captured using a mobile monocular camera.
@@
TABLE V: Scene decomposition results on real-world datasets.
@@
Dynamic scene editing. The object-centric representations acquired through DynaVol-S demonstrate the capability for seamless integration into scene editing workflows, eliminating the need for additional training. By manipulating the 4D voxel grid $\mathcal{V}^{\text{D}}$ learned in DynaVol-S, objects can be replaced, removed, duplicated, or added according to specific requirements. For example, we can swap the color between objects by swapping the corresponding latent codes assigned to each object. Moreover, the deformation field of individual objects can be replaced with user-defined trajectories (e.g., rotations and translations), enabling precise animation of the objects in the scene.
@@ 4.2 Novel View Synthesis
Synthetic scenes. We compare the performance of DynaVol-S for novel view synthesis with D-NeRF, DeVRF, and TineuVox. As DeVRF is trained with $V=60$ views at the initial timestamp and $4$ views for subsequent timestamps, we also train a DeVRF model under a data configuration similar to ours. Specifically, we only have access to one image per timestamp ($t\geq 1$) in a dynamically sampled view from the upper hemisphere. This model is termed as “DeVRF-Dyn”. As shown in Table II, DeVRF-Dyn yields a significant decline in performance compared to its standard baseline, primarily due to its heavy reliance on accurate initial scene understanding. In contrast, DynaVol-S achieves the best results across most scenes among all the compared models.
Real-world scenes. We also evaluate the performance of DynaVol-S in real-world scenes with cutting-edge NeRF-based methods, including NeurDiff, HyperNeRF, TineuVox, and $\text{D}^{2}$eRF. As shown in Table III, DynaVol-S outperforms the previous state-of-the-art method, HyperNeRF, by $10.8\%$ in PSNR and $8.3\%$ in SSIM on average. Furthermore, we also compare DynaVol-S with another neural rendering method that learns disentangled representations. The results indicate that our approach outperforms $\text{D}^{2}$eRF by $8.2\%$ in PSNR and $3.5\%$ in SSIM.
@@
Qualitative results. Fig. 3 showcases the rendered images at an arbitrary timestamp from a novel view. On the synthetic dataset, it shows that DynaVol-S captures 3D geometries and the motion patterns of different objects more accurately than the compared methods. In contrast, D-NeRF fails to model the complex motion in 3ObjRand and generates blurry results in 3ObjRealCpmx. DeVRF, on the other hand, also fails to accurately model the trajectories of the moving object with complex textures, as highlighted by the red box in 3ObjRealCmpx. The real-world examples on the right side in Fig. 3 illustrate that DynaVol-S generates more high-quality novel view images than the prior art, TineuVox.
@@ 4.3 Scene Decomposition
Synthetic scenes. To obtain quantitative results for 2D segmentation as well as the scene decomposition maps, we assign the casted rays in volume rendering to different latent codes according to each latent code’s contribution to the ray’s final color. Table IV provides a comparison between DynaVol-S with 2D/3D unsupervised object-centric decomposition methods (SAVi, uORF, and MovingParts) and a pretrained SAM. For a fair comparison with SAVi, which works on 2D video inputs, we implement a DynaVol-S model using consecutive images with a fixed camera view (termed as FixCam). For uORF and SAM, as they are primarily designed for static scenes, we preprocess the dynamic sequence into $T$ individual static scenes as the inputs of these models and evaluate their average performance on the whole sequence. For MovingParts, which generates hierarchical clustering results by iteratively merging the slots with the most similar motions, we select the outcome that achieves the highest mIOU as its final result. The FG-ARI and mIOU results in Table IV show that DynaVol-S outperform all of the compared models by large margins.
Real-world scenes. We further evaluate the scene decomposition results of DynaVol-S in real-world scenes against the state-of-the-art unsupervised video segmentation method, OCLR [15]. Notably, OCLR utilizes DINOv1 features [32] for test-time adaptation, referred to as OCLR (Test Adapt), to enhance temporal consistency. Additionally, we compare DynaVol-S with the open-world object segmentation approach, ClipSeg [16]. Since ClipSeg relies on text prompts, we manually adjust the prompts to tune the model. It is important to highlight that ClipSeg employs the CLIP model [33] and is further trained with true segmentation in a supervised manner, whereas our approach is entirely unsupervised. As demonstrated in Table V, DynaVol-S consistently achieves the best decomposition results, significantly outperforming all compared methods.
@@
TABLE VI: Ablation studies of the key components and the training stages in DynaVol-S.
@@
Qualitative results. In Fig. 4, we randomly select a timestamp on 6ObjFall and present the object-centric decomposition maps of SAVi, uORF, SAM, MovingParts, and DynaVol-S. We have the following three observations from the visualized examples. First, compared with the 2D models (SAVi and SAM), DynaVol-S can effectively handle severe occlusions between objects in 3D space. It shows the ability to infer the complete shape of the objects, as illustrated in the $1$st and $5$th latent codes. Second, compared with the 3D decomposition uORF, DynaVol-S can better segment the dynamic scene by leveraging explicitly meaningful spatiotemporal representations, while uORF only learns latent representations for each object. Furthermore, DynaVol-S is shown to adaptively work with redundant latent codes, in the sense that the pre-defined number of latent codes can be larger than the actual number of objects in the scene. In such cases, the additional latent codes learn to disentangle noise in visual observations or learn to not contribute significantly to image rendering, enhancing the model’s flexibility and robustness. In Fig. 5, we showcase the real-world scene decomposition results of DynaVol-S compared with OCLR and ClipSeg. Our findings reveal that DynaVol-S produces clearer 3D segmentation for each object, especially at object boundaries, while also maintaining temporal consistency across the video frames. In contrast, OCLR struggles to preserve temporal consistency. DynaVol-S also outperforms its our previous work, DynaVol, indicating the importance and effectiveness of incorporating semantic features. More real-world examples can be found in Fig. 6 and in the Appendix Fig. 11– 13.
TABLE VII: The analysis of the impact of the number of latent codes ($N$) on the performance of our approach.
# Latent Codes 3ObjRand 8ObjFall 3ObjRealCmpx
[…]
@@
TABLE VIII: Hyperparameter choice of the threshold value for splitting the foreground and background in object-centric voxel grids initialization stage, and the weight of per-point RGB loss ($\alpha_{p}$).

### v2

@@ 3.4 Training Procedure
DynaVol-S involves three stages in the training phase: First, a warmup stage that learns to obtain a basic understanding of the geometric, semantic, and dynamic priors. Second, an initialization stage for the voxelized occupancy probabilities ${\mathcal{V}^{\text{Occ}}}$, involving the connected components algorithm. Third, an end-to-end optimization stage that jointly optimizes $\{{\mathcal{V}^{\text{Opac}}},{\mathcal{V}^{\text{Color}}},{\mathcal{V}^{\text{Occ}}},\{\mathbf{z}_{n}\}_{n=1}^{N},f_{\psi}\}$ to refine the voxelized object occupancy.
@@
20:     ${\mathcal{V}^{\text{Occ}}}[n,:,:,:]=\text{one-hot}(\text{cl}_{k})$triangleright Generate 4D voxel grids
@@
22:  Rendering 2D masks $\widehat{P}(\mathbf{r})$ with ${\mathcal{V}^{\text{Occ}}}$
@@ 3.4.1 Warmup
To reduce the difficulty of learning the object-centric 4D voxel grids ${\mathcal{V}^{\text{Occ}}}$, we first use $\{\mathbf{I}_{t}\}_{t=1}^{T}$ and $\{\mathbf{F}_{t}\}_{t=1}^{T}$ to warm-up the density grids ${\mathcal{V}^{\text{Opac}}}$, the color-related feature grids ${\mathcal{V}^{\text{Color}}}$, the semantic feature grids ${\mathcal{V}^{\text{Sem}}}$, the slot features $\mathbf{S}$, and the dynamics module $f_{\psi}(\cdot)$. Additionally, we train an extra forward dynamics module $f^{\prime}_{\xi}(\cdot)$ and a non-compositional neural renderer $N_{\phi^{\prime}}(\cdot)$. The overall objective is defined as $\mathcal{L}_{\text{Warm}}=\mathcal{L}_{\text{Render}}+\alpha_{p}\mathcal{L}_{\text{Point}}+\alpha_{b}\mathcal{L}_{\text{Bg-Ent}}+\alpha_{c}\mathcal{L}_{\text{Cyc}}+\mathcal{L}_{\text{Recon}}+\alpha_{e}\mathcal{L}_{\text{Ent}}$. This framework divides its focus between scene reconstruction, covered by the first four terms, and the representation of semantic features, addressed by the last two terms. The hyperparameter values of scene reconstruction are adopted from prior literature [20]. At a specific time, we define
@@
where $\mathcal{L}_{\text{Render}}$ is the rendering loss $\mathcal{L}_{\text{Render}}$ between the predicted and observed pixel colors, $\mathcal{L}_{\text{Bg-Ent}}$ is the background entropy loss that encourages the renderer to concentrate on either foreground or background, and $\mathcal{L}_{\text{Point}}$ is the per-point RGB loss following DVGO [23]. $\mathcal{R}$ is the set of sampled rays in a batch, $P$ is the number of sampling points along ray $r$, and $\widehat{w}_{l}^{r}$ is the color contribution of the last sampling point obtained by $\widehat{w}^{r}_{l}=T_{P}(1-\exp(-\sigma_{P}\delta_{P}))$. We use the above loss functions as basic objective terms for neural rendering.
To enhance the coherence of the learned canonical-space dynamics, we train an additional module $f^{\prime}_{\xi}(\cdot)$ using the following cycle-consistency loss:
@@
where $x^{\prime}_{i}=x_{i}+f_{\psi}(x_{i},t),i\in[1,P]$. Notably, since both networks, $f^{\prime}_{\xi}(\cdot)$ and $f_{\psi}(\cdot)$, are jointly trained during the same stage, it raises the question of whether they can be correctly trained together. In Sec. 4.5, we provide further discussions and experimental analyses to address this concern.
@@
Furthermore, we incorporate the reconstruction loss $\mathcal{L}_{\text{Recon}}$ between the predicted and observed pixel semantic features, as well as the entropy loss $\mathcal{L}_{\text{Ent}}$ on the semantic probabilities to regularize them towards a one-hot distribution:
3.4.2 Object-Centric Voxel Grids Initialization
We initialize the 4D object-centric grids $\mathcal{V}^{\text{M}}$ with two steps: (1) the feature graph generation and connected components computation and (2) a post-processing step with Conditional Random Field (CRF). As illustrated in Alg. 1, the entire process takes as inputs the voxel position set $\{\mathbf{x}_{k}\}$ with a size of $N_{s}=N_{x}\times N_{y}\times N_{z}$, the voxel density set $\{\sigma_{k}\}$ from ${\mathcal{V}^{\text{Opac}}}$, the view-independent colors $\{\mathbf{c}_{k}^{\text{ind}}\}$ generated by $N_{\phi^{\prime}}$, and the corresponding semantic labels $\{y_{k}\}$ obtained by $\mathrm{argmax}_{N}\left(\mathcal{A}\right)$ in Eq. (5). We also consider the forward canonical-space transitions generated by $f^{\prime}_{\xi}(\cdot)$. Besides, we set five hyperparameters in advance, where $\delta_{\text{den}}$ represents the density value threshold, $\delta_{\text{pos}}$ represents the position distance threshold, $\delta_{\text{rgb}}$ represents the RGB distance threshold, $\delta_{\text{vel}}$ represents the velocity distance threshold, and $N$ represents the maximum number of objects.
Furthermore, as an optional post-processing step in DynaVol-S, we employ a Conditional Random Field (CRF) to refine $\mathcal{V}^{M}$, following the common practice in video object segmentation, this step is beneficial for mitigating potential noise introduced by the connected components algorithm, thereby smoothing the segmentation results. The CRF method leverages RGB information to refine 2D mask predictions. Although this step could potentially compromise the 4D consistency of segmentation results, we address this by using the refined masks as pseudo-labels in the subsequent stage. This allows us to finetune the 4D voxel grids, ${\mathcal{V}^{\text{Occ}}}$, thereby ensuring consistency in object-centric representations.
3.4.3 Multi-Grids Joint Optimization
In this stage, we finetune the dynamics module $f_{\psi}$, the geometric grids ${\mathcal{V}^{\text{Opac}}}$, the appearance grids ${\mathcal{V}^{\text{Color}}}$, and the object-centric grids ${\mathcal{V}^{\text{Occ}}}$ obtained in the first two stages. We finetune the parameters of $({\mathcal{V}^{\text{Opac}}},{\mathcal{V}^{\text{Color}}},{\mathcal{V}^{\text{Occ}}},\psi)$ and begin training ($(\{\mathbf{z}_{n}\}_{n=1}^{N},\phi)$ from scratch. Following an end-to-end training scheme, the dynamics module, object-centric latent codes, and compositional renderer collaboratively contribute to object-centric dynamic scene understanding. The loss function in this stage is defined as $\mathcal{L}_{\text{Dyn}}=\mathcal{L}_{\text{Render}}+\alpha_{p}\mathcal{L}_{\text{Point}}+\alpha_{b}\mathcal{L}_{\text{Bg-Ent}}$, complemented by a cross-entropy loss between the rendered 2D segmentations $\widehat{P}(\mathbf{r})$ from ${\mathcal{V}^{\text{Occ}}}$ and the pseudo-labels $\bar{P}(\mathbf{r})$ derived from the second stage: $\mathcal{L}_{\text{CE}}=-\sum_{n=1}^{N}\bar{P}(\mathbf{r})_{n}\log\widehat{P}(\mathbf{r})_{n}$.
4 Experiments
@@ 4.1 Experimental Setup
Datasets. We build the $8$ synthetic dynamic scenes in Table II using the Kubric simulator [25]. Each scene spans $60$ timestamps and contains different numbers of objects in various colors, shapes, and textures. The objects have diverse motion patterns and initial velocities. All images have dimensions of $512\times 512$ pixels. We also adopt real-world scenes from HyperNeRF [8] and $\text{D}^{2}$eRF [9], as shown in Table III and Table VI. For the synthetic scenes, we follow D-NeRF [17] to employ images collected at viewpoints randomly sampled on the upper hemisphere. In contrast, real-world scenes are captured using a mobile monocular camera.
@@
TABLE V: 3D scene decomposition results on the synthetic datasets in terms of 3D-ARI.
Method 3ObjFall 3ObjRand 3ObjMetal 3Fall+3Still 6ObjFall 8ObjFall 3ObjRealSimp 3ObjRealCmpx
MovingParts 27.02 37.94 48.24 25.52 27.89 42.27 46.12 42.82
DynaVol 87.46 42.10 66.51 77.55 88.21 80.24 62.25 63.00
DynaVol-S 90.02 94.37 87.44 84.29 88.20 81.37 91.00 92.88
@@
TABLE VI: Scene decomposition results on real-world datasets.
@@
Dynamic scene editing. The object-centric representations acquired through DynaVol-S demonstrate the capability for seamless integration into scene editing workflows, eliminating the need for additional training. By manipulating the 4D voxel grid ${\mathcal{V}^{\text{Opac}}}$ learned in DynaVol-S, objects can be replaced, removed, duplicated, or added according to specific requirements. For example, we can swap the color between objects by swapping the corresponding latent codes assigned to each object. Moreover, the deformation field of individual objects can be replaced with user-defined trajectories (e.g., rotations and translations), enabling precise animation of the objects in the scene.
@@ 4.2 Novel View Synthesis
We compare the performance of DynaVol-S for novel view synthesis with D-NeRF, DeVRF, and TineuVox. As DeVRF is trained with $V=60$ views at the initial timestamp and $4$ views for subsequent timestamps, we also train a DeVRF model under a data configuration similar to ours. Specifically, we only have access to one image per timestamp ($t\geq 1$) in a dynamically sampled view from the upper hemisphere. This model is termed as “DeVRF-Dyn”. As shown in Table II, DeVRF-Dyn yields a significant decline in performance compared to its standard baseline, primarily due to its heavy reliance on accurate initial scene understanding. In contrast, DynaVol-S achieves the best results across most scenes among all the compared models.
We also evaluate the performance of DynaVol-S in real-world scenes with cutting-edge NeRF-based methods, including NeurDiff, HyperNeRF, TineuVox, and $\text{D}^{2}$eRF. As shown in Table III, DynaVol-S outperforms the previous state-of-the-art method, HyperNeRF, by $10.8\%$ in PSNR and $8.3\%$ in SSIM on average. Furthermore, we also compare DynaVol-S with another neural rendering method that learns disentangled representations. The results indicate that our approach outperforms $\text{D}^{2}$eRF by $8.2\%$ in PSNR and $3.5\%$ in SSIM.
@@
Fig. 3 showcases the rendered images at an arbitrary timestamp from a novel view. On the synthetic dataset, it shows that DynaVol-S captures 3D geometries and the motion patterns of different objects more accurately than the compared methods. In contrast, D-NeRF fails to model the complex motion in 3ObjRand and generates blurry results in 3ObjRealCpmx. DeVRF, on the other hand, also fails to accurately model the trajectories of the moving object with complex textures, as highlighted by the red box in 3ObjRealCmpx. The real-world examples on the right side in Fig. 3 illustrate that DynaVol-S generates more high-quality novel view images than the prior art, TineuVox.
@@ 4.3 Scene Decomposition
To obtain quantitative results for 2D segmentation as well as the scene decomposition maps, we assign the casted rays in volume rendering to different latent codes according to each latent code’s contribution to the ray’s final color. Table IV provides a comparison between DynaVol-S with 2D/3D unsupervised object-centric decomposition methods (SAVi, uORF, and MovingParts) and a pretrained SAM. For a fair comparison with SAVi, which works on 2D video inputs, we implement a DynaVol-S model using consecutive images with a fixed camera view (termed as FixCam). For uORF and SAM, as they are primarily designed for static scenes, we preprocess the dynamic sequence into $T$ individual static scenes as the inputs of these models and evaluate their average performance on the whole sequence. For MovingParts, which generates hierarchical clustering results by iteratively merging the slots with the most similar motions, we select the outcome that achieves the highest mIOU as its final result. The FG-ARI and mIOU results in Table IV show that DynaVol-S outperform all of the compared models by large margins.
Furthermore, we assess scene decomposition in 3D space. Specifically, we first voxelize the ground truth mesh for each object in the synthetic scenes to obtain their surface points. These surface points are then used to query the corresponding predicted labels in $\mathcal{V}^{\text{Occ}}$. The predicted labels are compared to the ground truth by calculating the 3D-ARI for each scene, as presented in Table V. The results demonstrate that DynaVol-S consistently outperforms other baselines across all synthetic datasets.
In Fig. 4, we randomly select a timestamp on 6ObjFall and present the object-centric decomposition maps of SAVi, uORF, SAM, MovingParts, and DynaVol-S. We have the following three observations from the visualized examples. First, compared with the 2D models (SAVi and SAM), DynaVol-S can effectively handle severe occlusions between objects in 3D space. It shows the ability to infer the complete shape of the objects, as illustrated in the $1$st and $5$th latent codes. Second, compared with the 3D decomposition uORF, DynaVol-S can better segment the dynamic scene by leveraging explicitly meaningful spatiotemporal representations, while uORF only learns latent representations for each object. Furthermore, DynaVol-S is shown to adaptively work with redundant latent codes, in the sense that the pre-defined number of latent codes can be larger than the actual number of objects in the scene. In such cases, the additional latent codes learn to disentangle noise in visual observations or learn to not contribute significantly to image rendering, enhancing the model’s flexibility and robustness. We further provide mesh visualization in Sec. 4.5 to demonstrate the ability of our approach for disentanglement representation learning.
We further evaluate the scene decomposition results of DynaVol-S in real-world scenes against the state-of-the-art unsupervised video segmentation method, OCLR [15]. Notably, OCLR utilizes DINOv1 features [32] for test-time adaptation, referred to as OCLR (Test Adapt), to enhance temporal consistency. Additionally, we compare DynaVol-S with the open-world object segmentation approach, ClipSeg [16]. Since ClipSeg relies on text prompts, we manually adjust the prompts to tune the model. It is important to highlight that ClipSeg employs the CLIP model [33] and is further trained with true segmentation in a supervised manner, whereas our approach is entirely unsupervised. As demonstrated in Table VI, DynaVol-S consistently achieves the best decomposition results, significantly outperforming all compared methods.
@@
TABLE VII: Ablation studies of the key components and the training stages in DynaVol-S. Additionally, we showcase the effect of using alternative designs for volume slot attention in Fig. 9.
@@
In Fig. 5, we showcase the real-world scene decomposition results of DynaVol-S compared with OCLR and ClipSeg. Our findings reveal that DynaVol-S produces clearer 3D segmentation for each object, especially at object boundaries, while also maintaining temporal consistency across the video frames. In contrast, OCLR struggles to preserve temporal consistency. DynaVol-S also outperforms our previous work, DynaVol, indicating the importance and effectiveness of incorporating semantic features.
Can DynaVol-S handle scenarios where objects leave or new objects enter the scene? The answer is yes. For example, in the left case in Fig. 5, the inner part of the Chicken toy (the yellow one) repeatedly leaves and reappears throughout the sequence, yet DynaVol-S successfully generates consistent labels. This robustness is primarily due to DynaVol-S performing segmentation in the canonical space, where all objects appearing in the sequence are represented, regardless of whether they are visible in specific frames. More real-world examples with severe occlusions can be found in Fig. 6 and supplementary materials.
TABLE VIII: The analysis of the impact of the number of latent codes ($N$) on the performance of our approach.
$N$ 3ObjRand 8ObjFall 3ObjRealCmpx
PSNR mIOU PSNR mIOU PSNR mIOU
@@
4.4 Dynamic Scene Editing
After the training period, the object-centric voxel representations learned by DynaVol-S can be readily used in downstream tasks such as scene editing without the need for additional model tuning. DynaVol-S allows for easy manipulation of the observed scene by directly modifying the object occupancy values within the voxel grids or switching the learned deformation function to a pre-defined one. This flexibility empowers users to make various scene edits and modifications. For instance, in the first example in Fig. 7, we remove the hand that is pinching the toys. In the second example, we modify the dynamics of the shoe from falling to rotating. More showcases are included in the supplementary materials.
4.5 Model Analyses
Ablation studies. Table VII presents the ablation study results, highlighting the effectiveness of the key component of DynaVol-S. First, we observe that object-centric voxelization enables effective scene decomposition and improves the rendering results. Subsequently, we can find that the forward deformation network $f_{\xi}^{\prime}(\cdot)$ is crucial to the final rendering and decomposition results. Furthermore, in the absence of the multi-grids joint optimization stage, the performance of DynaVol-S significantly degrades, highlighting the importance of refining the object-centric voxel representation with a latent-code-based renderer. Additionally, we conduct an ablation study that performs multi-grids joint optimization from scratch with randomly initialized ${\mathcal{V}^{\text{Occ}}}$ and $f_{\phi}(\cdot)$. The results clearly show that excluding the warmup stage substantially impacts the final performance, especially for the scene decomposition results.
The number of object-centric latent codes. First, we examine the impact of the number of latent codes ($N$), which indicates the maximum number of objects in the scene. According to Table VIII, the model achieves comparable results as long as $N$ exceeds the number of objects in the scene. The presence of redundant latent codes has only a minor influence on both rendering and decomposition results. Additionally, we showcase the per-object rendering results in supplementary materials, demonstrating DynaVol-S’s robustness to variations in the number of latent codes. Our findings reveal that redundant latent codes typically remain unassigned, effectively addressing the potential issue of over-segmentation.
TABLE IX: Hyperparameter analyses of (1) the threshold value for splitting the foreground and background in the object-centric voxel grids initialization stage, and (2) the weight of the per-point RGB loss ($\alpha_{p}$). We use $\delta_{\text{den}}=1,\alpha_{p}=0.01$ for the final model.